OpenStudy (anonymous):

Which conjunction or disjunction is equivalent to the open sentence |a – 9| ≤ 4. A. a – 9 ≥ 4 and a – 9 ≤ –4 B. a – 9 ≤ 4 or a – 9 ≥ –4 C. a – 9 ≤ 4 and a – 9 ≤ –4 D. a – 9 ≤ 4 and a – 9 ≥ –4 I have 4 other questions I need help with, most/all of them are screenshots though, I will fan and medal if you can help me with all of them.

OpenStudy (anonymous):

Which conjunction or disjunction equivalent to the open sentence 4 –3|n + 2| ≥ 1? A. n + 2 ≥ 1 and n + 2 ≤ –1 B. n + 2 ≤ 1 and n + 2 ≥ –1 C. n + 2 ≤ 1 and n − 2 ≥ −1 D. n + 2 ≥ 1 or n – 2 ≤ 1

OpenStudy (anonymous):

3. Solve the equation. |k + 9| = 4 A. {–13, –5} B. {–5, 5} C. {–5, 13} D. {all real numbers greater than or equal to –13 and less than or equal to –5}

OpenStudy (anonymous):

4. Solve the inequality. 3 + |t + 4| < 10 A. –3 < t < 11 B. t > 3 or t < –11 C. –11 < t < 3 <--- I think its this one. D. t < 3 and t > 11

OpenStudy (anonymous):

This is the last question.

OpenStudy (anonymous):

@dan815 @nincompoop @hartnn @triciaal can any of yall help me. I will do anything to get help from someone. I'm in a rush cause i've been trying to figure these out for like 3 hours now.

OpenStudy (anonymous):

I think maybe the first one A

OpenStudy (anonymous):

Wait no B

hartnn (hartnn):

yes, B for 1st is correct :) if you tell me how, I'll help you with 2nd question.

OpenStudy (anonymous):

And I think the 3rd one is C.

hartnn (hartnn):

so we are at 3rd one? when |a| = b, either a= b or a =-b so what can you say when |k+9| = 4 ??

OpenStudy (anonymous):

Well it took me longer to answer the 2nd one but I think its D.

OpenStudy (anonymous):

I meant i think the 2nd one is D

OpenStudy (anonymous):

Wait no I think the 3rd one is A

hartnn (hartnn):

lets skip the guessing work and actually solve the problem :) 3rd one : when |a| = b, either a= b or a =-b when |k+9| = 4 k+9 = 4 or k+9 = -4 can you solve for k in k+9 = 4? ?

OpenStudy (anonymous):

I know -5 + 9 =4 , and -13 + 9 = -4

hartnn (hartnn):

right, so thats how the 3rd one is A :)

hartnn (hartnn):

now 2nd one when |a| \(\ge \)b , either a \(\ge \)b or a \(\le \) -b

hartnn (hartnn):

4 –3|n + 2| ≥ 1 try to isolate |n+2| from that first

OpenStudy (anonymous):

Is it -1

hartnn (hartnn):

how did you get -1? very close because i got +1 |n+2| \(\le \) +1

hartnn (hartnn):

|n+2| ≤ +1 means n+2 ≤ +1 or n+2 >= -1 which is B

OpenStudy (anonymous):

Am I correct on number 4?

OpenStudy (anonymous):

And for 5 i think its C.

hartnn (hartnn):

4th is correct! :) and 5th is indeed C :)

OpenStudy (anonymous):

I went ahead and submitted it and made an 80! They only one that was wrong was number 1, the correct answer was D.

hartnn (hartnn):

oh so its an "and" instead of "or" :P and \(\Huge \mathcal{\text{Welcome }\ddot\smile} \)